Write the equation of the line parallel to 2y-8x=-10 that goes through the point (0,-5).
Answer by Simnepi(216)

The form y = mx + b gives the equation of a straight line where m is the slope of the line(the gradient) and b is the intercept of the line and the y axis.
All lines with the same gradient (i.e. the same number in front of the x) are parallel.
First, let's rearrange your line equation into the correct form of y = mx + b
2y-8x=-10
Move the x term across the equals sign, changing its sign as we do so
2y = 8x - 10
Now divide by 2 (through the whole equation) to get rid of the 2y and just leave y
y = 4x - 5
Now we can see that the line we want has a gradient of 4.
it also intercepts the y axis when y = -5
The line we want is therefore
y = 4x - 5
It is the same line as the one given!